OpenStudy (anonymous):

Evaluate the following limit, if it exists. Give exact value lim (2-5/x + 7/x^2) x->-infinity

OpenStudy (anonymous):

\[\lim_{x \rightarrow -\infty} \left( 2-\left(\begin{matrix}5 \\ x\end{matrix}\right) + \frac{ 7 }{ x^{2} }\right)\]

OpenStudy (anonymous):

2

OpenStudy (turingtest):

proof please

OpenStudy (anonymous):

Ok...could you tell me how you got that answer?

OpenStudy (anonymous):

It is open secret I don't need to poof it.

OpenStudy (anonymous):

Ok well thanks anyways. I'm just trying to understand how to do this question, not looking for just the answer

OpenStudy (anonymous):

if you sub in -infinity for x, the last 2 terms are insignificant and can be ignored.

OpenStudy (turingtest):

the deal is that as x->infty, 1/x->0 hence any finite number divided by \(\pm\infty\) is zero so \[\lim_{x\to\infty}\frac5x=\frac5\infty=0\]and\[\lim_{x\to\infty}\frac7{x^2}=\frac7\infty=0\]so all that's left is the 2

OpenStudy (turingtest):

same for negative infinity...

OpenStudy (anonymous):

oh ok, thank u petewe. So if it was x->2 I would just substitute that number in for x?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

oh wow, that's easy enough. Thank you both petewe and Turning Test!

OpenStudy (turingtest):

happy to help :)